# Atom of Thoughts for Markov LLM Test-Time Scaling

### v1

@@ Abstract
Large Language Models (LLMs) achieve superior performance through training-time scaling, and test-time scaling further enhances their capabilities by conducting effective reasoning during inference. However, as the scale of reasoning increases, existing test-time scaling methods suffer from accumulated historical information, which not only wastes computational resources but also interferes with effective reasoning. To address this issue, we observe that complex reasoning progress is often achieved by solving a sequence of independent subquestions, each being self-contained and verifiable. These subquestions are essentially atomic questions, relying primarily on their current state rather than accumulated history, similar to the memoryless transitions in a Markov process. Based on this observation, we propose Atom of Thoughts (AoT), where each state transition in the reasoning process consists of decomposing the current question into a dependency-based directed acyclic graph and contracting its subquestions, forming a new atomic question state. This iterative decomposition-contraction process continues until reaching directly solvable atomic questions, naturally realizing Markov transitions between question states. Furthermore, these atomic questions can be seamlessly integrated into existing test-time scaling methods, enabling AoT to serve as a plug-in enhancement for improving reasoning capabilities. Experiments across six benchmarks demonstrate the effectiveness of AoT both as a standalone framework and a plug-in enhancement. Notably, on HotpotQA, when applied to gpt-4o-mini, AoT achieves an 80.6% F1 score, surpassing o3-mini by 3.4% and DeepSeek-R1 by 10.6%. The code will be available at https://github.com/qixucen/atom.
@@ 1 Introduction
Large Language Models (LLMs) demonstrate significant scaling effects, with their capabilities showing predictable improvements as model parameters and training data increase, leading to enhanced performance across diverse domains Kaplan et al. (2020). While this scaling law faces bottlenecks in high-quality data availability, test-time scaling offers an alternative solution by forcing LLMs to engage in effective logical reasoning during inference to improve performance on diverse tasks Snell et al. (2024); Muennighoff et al. (2025); Hou et al. (2025); Zhang et al. (2024a).
However, existing test-time scaling methods excessively maintain historical information during reasoning, as they rely heavily on complex structural dependencies throughout the reasoning process. Chain-based methods must preserve the entire reasoning history to generate each subsequent step Wei et al. (2022); Zhang et al. (2023), while tree-based approaches require tracking both ancestor and sibling relationships for branch selection Yao et al. (2023); Zhou et al. (2024); Ding et al. (2024). Graph-based structures further compound these challenges through arbitrary node dependencies. As the scale of reasoning increases, the accumulation of historical dependencies not only wastes substantial computational resources but also interferes with the model’s ability to reason effectively, as illustrated in Figure 1.
@@
  • Plug-In Enhancement. The atomic questions derived by AoT can be directly integrated into existing test-time scaling methods Bi et al. (2024); Wang et al. (2023b), enhancing both their performance and cost efficiency.
@@ 2.1 Reasoning Framework
Chain-of-Thought Wei et al. (2022) prompting has emerged as a fundamental technique for enhancing LLMs’ reasoning. Decomposition methods like Least-to-Most Zhou et al. (2023) and Plan-and-Solve Wang et al. (2023a) prompting parse complex problems into sequential subtasks. Iterative optimization approaches like Self-Refine Madaan et al. (2023), Step-Back Zheng et al. (2024) prompting and Progressive-Hint Prompting Zheng et al. (2023) refine solutions through cyclic feedback or abstraction. Multi-path aggregation techniques like Self-Consistency CoT Wang et al. (2023b) and LLM-Blender Jiang et al. (2023) further improve reasoning reliability by multi-trajectory consensus.
More sophisticated frameworks structure the representation of reasoning space through dedicated formalisms: Tree of Thoughts  Yao et al. (2023) enables systematic exploration of multiple reasoning paths, while Graph of Thoughts Besta et al. (2024) represents reasoning processes as dynamic graphs with backtracking mechanisms. Addressing fundamental limitations in resampling-based paradigms, Thought Space Explorer Zhang and Liu (2024) strategically explores under-sampled regions of the solution space. These frameworks serve as universal augmentation of LLMs reasoning, enhancing their capacity across various domains, with their principles being widely adopted in agentic workflows for code generation, question answering, and data science applications Hong et al. (2024b); Zhang et al. (2024a); Hong et al. (2024a); Zhang et al. (2025); Xiang et al. (2025); Zhang et al. (2024b).
@@ 2.2 Test-time Scaling
Test-time scaling approaches have demonstrated the value of extended computation during inference. Supervised fine-tuning on long chain-of-thought traces has proven effective at enhancing models’ capabilities to conduct extended reasoning Yeo et al. (2025); Yao et al. (2025). Building on this foundation, reinforcement learning methods have enabled models to automatically learn optimal inference expansion strategies, allowing for adaptive scaling of the reasoning process Kimi et al. (2025); Zeng et al. (2025); DeepSeek-AI (2025). Framework-based approaches have further expanded these capabilities by extending inference through external systems, incorporating techniques like verification, budget forcing, and ensemble methods Zhang et al. (2024a); Saad-Falcon et al. (2024); Chen et al. (2024). These complementary approaches demonstrate how strategic use of additional computation during inference through learned behaviors, automated scaling, and system-level interventions can substantially enhance model performance.
However, these approaches universally maintain extensive historical information throughout the reasoning process, leading to computational inefficiency and potential interference with effective reasoning. In contrast, AoT introduces a Markovian perspective that eliminates the need for historical dependency tracking, enabling more efficient resource allocation while maintaining compatibility with existing test-time scaling methods.

### v2

@@ Abstract
Large Language Models (LLMs) achieve superior performance through training-time scaling, and test-time scaling further enhances their capabilities by conducting effective reasoning during inference. However, as the scale of reasoning increases, existing test-time scaling methods suffer from accumulated historical information, which not only wastes computational resources but also interferes with effective reasoning. To address this issue, we observe that complex reasoning can be achieved by solving a series of independent and self-contained subquestions. These subquestions are essentially atomic questions, exhibiting the memoryless property similar to Markov processes. Based on this observation, we propose Atom of Thoughts (AoT), where each state transition consists of decomposing the current question into a dependency-based directed acyclic graph and contracting its subquestions, forming a simplified question that maintains answer equivalence with the original problem. This answer preservation enables the iterative decomposition-contraction process to naturally form a meaningful Markov reasoning process. Furthermore, these atomic states can be seamlessly integrated into existing test-time scaling methods, enabling AoT to serve as a plug-in enhancement for improving reasoning capabilities. Experiments across six benchmarks demonstrate the effectiveness of AoT both as a standalone framework and a plug-in enhancement. Notably, on HotpotQA, when applied to gpt-4o-mini, AoT achieves an 80.6% F1 score, surpassing o3-mini by 3.4% and DeepSeek-R1 by 10.6%. The code is available at https://github.com/qixucen/atom.
@@ 1 Introduction
Large Language Models (LLMs) demonstrate significant scaling effects, with their capabilities showing predictable improvements as model parameters and training data increase, leading to enhanced performance across diverse domains Kaplan et al. (2020). While this scaling law faces bottlenecks in high-quality data availability Villalobos et al. (2024), test-time scaling offers an alternative solution by using more test-time computation to improve performance on diverse tasks Snell et al. (2024); Muennighoff et al. (2025); Hou et al. (2025).
However, existing test-time scaling methods excessively maintain historical information during reasoning, as they rely heavily on complex structural dependencies throughout the reasoning process. Chain-based methods must preserve the entire reasoning history to generate each subsequent step Wei et al. (2022); Zhang et al. (2023), while tree-based approaches require tracking both ancestor and sibling relationships for branch selection Yao et al. (2023); Zhou et al. (2024a); Ding et al. (2024). Graph-based structures further compound these challenges through arbitrary node dependencies Besta et al. (2024); Zhang et al. (2024c). As the scale of reasoning increases, the accumulation of historical dependencies not only wastes substantial computational resources but also interferes with the model’s ability to reason effectively, as illustrated in Figure 1.
@@
  • Plug-In Enhancement. The atomic question derived by AoT can be directly integrated into existing test-time scaling methods Bi et al. (2024); Wang et al. (2023b), enhancing both their performance and cost efficiency.
@@ 2.1 Reasoning Framework
Inspired by cognitive operations Bechtel et al. (2001) in human reasoning—such as step-by-step decomposition Wei et al. (2022); Zhou et al. (2023); Wang et al. (2023a); Hao et al. (2024), reflective refinement Madaan et al. (2023); Zheng et al. (2024, 2023), and aggregation ensemble Wang et al. (2023b); Jiang et al. (2023); Yao et al. (2025b)—various prompting strategies have been widely applied to enhance the reasoning capabilities of large language models (LLMs). Reasoning frameworks leverage structured assumptions, such as chains, graphs, and trees Yao et al. (2023); Besta et al. (2024); Zhang and Liu (2024), to organize these strategies, thereby effectively exploring and modeling the reasoning space. For instance, chain-based approaches decompose complex problems into sequential subtasks Wei et al. (2022); Zhou et al. (2023); Wang et al. (2023a), while tree- and graph-based formalisms enable systematic exploration of multiple reasoning paths with dynamic adaptability Yao et al. (2023); Besta et al. (2024). Such frameworks augment LLMs’ reasoning across domains, integrating principles that enhance performance in agentic workflows like code generation, question answering, and data science applications Hong et al. (2024b, a); Zhang et al. (2025, 2024b). By structuring the reasoning process, these methods provide a robust foundation for LLMs to tackle intricate tasks with improved coherence and reliability.
Compared to the ongoing efforts to design increasingly complex structures, several attempts have explored a Markov reasoning process or atomic reasoning steps. However, these approaches often suffer from a lack of generalizability and true atomic design Xin et al. (2024); Xiang et al. (2024); Zhou et al. (2024b); Xiang et al. (2025). For example, the definition of states and transitions remains task-specific and disrupts potential parallelism Hao et al. (2023); Yang et al. (2024); Zekri et al. (2024). In contrast, our approach utilizes a Directed Acyclic Graph (DAG) composed of subproblem nodes and dependency edges to accommodate a wide variety of tasks. By decomposing and contracting the DAG, our method isolates independent subproblems while preserving support for parallelism. Furthermore, our approach defines atomic units as questions that trigger reasoning, aligning naturally with real reasoning scenarios and effectively separating information from historical dependencies.
@@ 2.2 Test-time Scaling
Test-time scaling approaches have demonstrated the value of extended computation during inference. Framework-based approaches enhance LLM reasoning performance by extending inference through external systems, incorporating techniques like cognitive operations and external tool invocation Zhang et al. (2024a); Saad-Falcon et al. (2024); Chen et al. (2024). However, these methods do not improve the inherent reasoning capabilities of LLMs at the model level. To address this gap, supervised fine-tuning on long chain-of-thought trajectoris has proven effective at enhancing models’ capabilities to conduct extended reasoning Ye et al. (2025); Chang et al. (2025); Yao et al. (2025a). Building on this foundation, reinforcement learning methods have enabled models to automatically learn to extend chains of thought, with the potential to emerge some autonomous cognitive operations Kimi et al. (2025); Zeng et al. (2025); DeepSeek-AI (2025); Yu et al. (2025).
However, these approaches universally retain extensive historical information throughout the reasoning process, resulting in computational inefficiency and potential interference with effective reasoning. In contrast, AoT introduces an approach that integrates a Markovian perspective with an atomized perspective, eliminating the need for historical dependency tracking by decomposing the reasoning process into independent atomic reasoning units. This ingenious unification of the two properties enables AoT to concentrate computational resources on the current state, thereby achieving more efficient resource allocation, while the atomic state definition provides plug-in compatibility with existing test-time scaling methods.